The decomposition of acetaldehyde in gas phase was studied at 729k using the initial concentration of 7.77×10^-3M and 3.64×10^-3M. The half life periods were determined to be 295s and 515s respectively. Determine the order of the reaction.

1
Expert's answer
2022-04-05T14:16:03-0400

t1/2(1) = 295 s;

t1/2(2) = 515 s;

C0(1) = 0.00777 M;

C0(2) = 0.00364 M;

Order №1:

t1/2 = ln2/k

k1 = ln2/t1/2(1) = 0.693/295 = 0.002;

k2 = ln2/t1/2(2) = 0.693/515 = 0.001;

Order №2:

t1/2 = 1/k * 1/C0;

k1 = 1/(t1/2(1) * C0(1)) = 1/(295 * 0.00777) = 0.4;

k2 = 1/(t1/2(2) * C0(2)) = 1/(515 * 0.00364) = 0.5;

Order №3:

t1/2 = 1/k * 3/(2 * C02);

k1 = 3/(2 * C0(1)2 * t1/2(1)) = 3/(2 * 0.007772 * 295) = 84.2;

k2 = 3/(2 * C0(2)2 * t1/2(2)) = 3/(2 * 0.003642 * 515) = 219.;


Therefore: for the reaction determine the second order, k1 = k2